OpenStudy (anonymous):

Solve: 8x − 10 + 18 = 24

OpenStudy (anonymous):

2

OpenStudy (ash2326):

We have \[8x-10+18=24\] First step combine these two \[8x\underline{-10+18}=24\] @terrell352 what would you get now?

OpenStudy (anonymous):

wouldnt yu add 10 to 24 an subtract 18 from 34 an get16 an divide 16 by 8 an get 2

OpenStudy (ash2326):

That's also a way, good work :D But it's always nice to know more than one way of doing something :D

OpenStudy (anonymous):

oh ok what would be another way of doing it

OpenStudy (anonymous):

an thankz

OpenStudy (ash2326):

combine the terms which I underlined

OpenStudy (anonymous):

then you 8x+8=24

OpenStudy (ash2326):

Yeah, good Now subtract 8 from both sides

OpenStudy (anonymous):

16

OpenStudy (ash2326):

I mean what do you get, ?

OpenStudy (anonymous):

2

OpenStudy (ash2326):

good, did you understand?

OpenStudy (anonymous):

YEA more than the last one

OpenStudy (ash2326):

good;D